OpenStudy (anonymous):

Solve. 7x − 2(3x + 4) =15

OpenStudy (anonymous):

7x-6x-8=15 x=23

OpenStudy (anonymous):

SHOW WORK MADAFIGLIO

OpenStudy (anonymous):

which part don't you understand?

OpenStudy (anonymous):

JUST SHOW THE WORK THATS ALL IM ASKING

OpenStudy (anonymous):

ok, expand the bracket, and that will get you -6x-8 then you add 8 to both sides so that you have x on the left, and integers on the right. my working is there.

OpenStudy (anonymous):

7x-6x-8=15 x-8=15 +8=+8 x=23